OpenStudy (anonymous):

Realted Rates Problem help! A girl flies a kite at a height of 200 ft, the wind carrying the kite horizontally away from her at a rate of 4 ft/sec. How fast must she let out the string when the kite is 428 ft away from her?

OpenStudy (anonymous):



OpenStudy (anonymous):

another pythagoras \[200^2+x^2=y^2\] \[2xx'=2yy'\] \[x'=4\] so \[4x=yy'\] and you want \(y'\) when \(y=428\)

OpenStudy (anonymous):

Nice work satellite. @payal0912: also bear in mind that when you are taking the derivative of x and y, they are both functions of time. Related rates always are implicitly defined as functions of time.

OpenStudy (anonymous):

so, y'=dy/dt

OpenStudy (anonymous):

take care

OpenStudy (anonymous):

I'm still confused on where to get the x value from

OpenStudy (anonymous):

hi paya are you in calculus class or in algebra or trig ?

OpenStudy (anonymous):

calculus

OpenStudy (anonymous):

there are two unknowns in this equation. i tried using the pythagorean to get x but my answer is incorrect. I got \[\sqrt{143184}\] as my x value using y=428 and z=200 for z^2 + x^2 = y^2. but then using that x value in the differentiation, I got an incorrect answer